The ratio of the measure of a supplement of an angle to that of the complement of the angle is 5:2, what is the measure of the angle?
Answer by anjulasahay(30)

The ratio of the measure of a supplement of an angle to that of the complement of the angle is 5:2, what is the measure of the angle?
ans:
let A be the ang to be find out
hence according to the definition the supplement and complement of the an angle and the question give
(180-A)/(90-A)=5/2
or 2*(180 - A ) = 5 * (90 - A)
0r 360 - 2A = 450 - 5A
OR 5A-2A = 450-360
OR 3A = 90
OR A = 30
NOTE :SUPPLEMENT OF THE ANGLE IS THE ANGLE WHICH IS WHEN ADDED IN THE GIVEN ANGLE GIVES A TOTAL OF 180.
COMPLEMENT OF THE ANGLE IS THE ANGLE WHICH IS WHEN ADDED IN THE GIVEN ANGLE GIVES A TOTAL OF 90.